### v1

# MLOps with Microservices:A Case Study on the Maritime Domain

Abstract
This case study describes challenges and lessons learned on building Ocean Guard: a Machine Learning–Enabled System (MLES) for anomaly detection in the maritime domain. First, the paper presents the system’s specification, and architecture. Ocean Guard was designed with a microservices’ architecture to enable multiple teams to work on the project in parallel. Then, the paper discusses how the developers adapted contract-based design to MLOps for achieving that goal. As a MLES, Ocean Guard employs code, model, and data contracts to establish guidelines between its services. This case study hopes to inspire software engineers, machine learning engineers, and data scientists to leverage similar approaches for their systems.
Keywords:
Microservices MLOps Software Architecture Machine Learning Enabled Systems Maritime Domain Case Study
1 Introduction
Maritime transportation is the cornerstone of global trade and commerce, facilitating the movement of goods and commodities across vast distances. The maritime domain represents one of the most dynamic and complex operational environments in global logistics and transportation. It is characterized by continuous movement across international boundaries, varying regulatory frameworks, and vulnerability to unpredictable weather conditions. As such, maritime operations generate a large amount of heterogeneous data from multiple sources.
The digital transformation of the maritime industry has accelerated in recent years, driven by economic pressures, environmental regulations, and safety concerns [3]. According to the International Maritime Organization (IMO), ships now function as floating data centers, equipped with hundreds of sensors. Improving maritime navigation and communication by facilitating the implementation of safe data sharing between vessels and between vessels and shore-based systems through automation [1]. This complex web of information flows across organizational boundaries between shipping companies, port authorities, maritime safety agencies, and commercial partners [8].
This rich data ecosystem presents opportunities and challenges, including the potential for predictive analytics, anomaly detection, and route optimization using Machine Learning (ML) [2]. However, extracting actionable insights requires systems capable of ingesting, processing, and analyzing different data streams at scale while maintaining operational reliability in an environment where connectivity and conditions rapidly change [17]. Machine Learning Operations (MLOps) refers to a collection of best practices, sets of concepts, and development culture that aims to achieve that goal [10].
Since 2015, the microservices architectural style have emerged as a promising approach to address challenges of modularity, scalability, and resilience by decomposing complex systems into smaller, independently deployable services with well-defined interfaces [14, 16]. This approach can be particularly suited to the maritime domain’s nature, where specialized components can be designed to handle different maritime data sources and analytical needs.
The novel field of Software Engineering for Artificial Intelligence (SE4AI) studies the adaptation of Software Engineering techniques for the creation of ML-Enabled Systems (MLES) [9]. Since 2022, different reference architectures have been proposed for them [11, 6, 10]. Fundamentally, they describe these systems as containing multiple components, which can be divided across different services.
This case study describes the development of Ocean Guard: a ML-Enabled System (MLES) for anomaly detection in the maritime domain. Section 2 details the functional and non-functional requirements considered for the system. Section 3 gives an overview of the system architecture, designed with a microservices architecture for MLOps. Section 4 zooms in in the Ocean Guard API, designed with a hexagonal architecture with Domain Driven Design (DDD). Section 5 shows the different teams involved in the development of the tool. Section 6 describes how the developers adapted contract-based design to MLES for developing different components in parallel. Section 7 lists the main challenges during the development of the Ocean Guard tool, according to the experience of the authors. Section 8 then summarizes the key lessons learned, related to applying development techniques from microservices on ML-enabled system development. Finally, section 9 lists the main contributions of paper and future work.
2 System Specification
This section describes the specification of the Ocean Guard tool, which was defined from conversations with stakeholders at the beginning of the MARIT-D project. The specification is divided into two parts: functional and non-functional requirements [15].
2.1 Functional Requirements
Functional Requirements, document the functionalities that directly affect the user of a system [15]. There are two actors related to the Ocean Guard tool. The following descriptions detail the main functionalities associated with them.
An Investigator is a human user who uses the tool to obtain clues about anomalous behavior at sea. They must be able to:
  1. I1.
    see geolocations of marine objects in a map,
  2. I2.
    filter geolocations by area of interest, date and time,
  3. I3.
    discern different types of marine objects (vessels, structures, or unidentified),
  4. I4.
    retrieve geolocations gathered from different data sources,
  5. I5.
    check details (metadata) associated with a given marine object,
  6. I6.
    highlight the trajectory of a marine object,
  7. I7.
    see anomalies identified by the tool in a map,
  8. I8.
    filter anomalies by area of interest, date and time, and
  9. I9.
    inspect why an anomaly was considered so by the tool.
The Anomaly Detector is a component of the tool responsible for detecting anomalous behavior of marine objects. It must be able to:
  1. A1.
    detect anomalies related to a marine object,
  2. A2.
    list anomalies by area of interest, date and time, and
  3. A3.
    explain why an anomaly is considered so.
2.2 Non-Functional Requirements
Non-Functional Requirements, also known as Architectural Requirements, describe a set of nonstructural attributes that are important for the success of a system [15]. There are four main non-functional requirements proposed for the Ocean Guard tool. The following descriptions detail the main characteristics the system should have to attend its users.
Compliance.
The Ocean Guard tool should be compliant with the latest European laws, including the General Data Protection Regulation (GDPR), and the Artificial Intelligence Act (AI Act).
Explainability.
The Ocean Guard tool should be able to explain why it considers a marine object to be anomalous. For algorithm-based anomaly detection, this means clarifying which decisions were taken until the anomaly was identified. For AI-based anomaly detection, this means using interpretable models (such as tree-based models) or implementing black-box explainability techniques (such as SHAP or LIME).
Resiliency.
The Ocean Guard tool should be able to handle the expected amount of data. This is important because the system may process high-volume and high-velocity data sources, such as those providing tracking data.
Compatibility.
The Ocean Guard tool should be able to interoperate with other existing tools, extending the Investigator’s current existing capabilities. As such, the tool should follow EU standards that enable quick integration with other ecosystems.
3 System Architecture
This section presents the architecture of the Ocean Guard tool, designed to implement the requirements presented at section 2. Figure 1 showcases the components planned for Ocean Guard, following the reference architecture representation proposed by Ferreira et al. [6].
Figure 1: System Architecture of the Ocean Guard tool This architecture follows the same notation introduced by Ferreira et al. [6]. Rectangles represent applications or services, which execute continuously. Stacked rectangles represent pipelines, which execute a task on demand. Lastly, cylinders represent data storage, which may be databases of any type. Components are connected by arrows. Black arrows with a hollow tip illustrate the execution flow. They start and end in a component. Labeled arrows represent the trigger that starts a workflow, whereas unlabeled arrows represent the continuation of an existing workflow. Colored arrows with a filled tip illustrate the data flow. They appear in two types: solid arrows going to and from a data storage represent write and read operations, respectively; dotted arrows represent a sync or async request-response communication between components. Components are colored according to the data they produce: raw data  , source code  , executable artifacts  , ML-specific data  , ML models  , ML training metadata  , ML model predictions  , and ML model metrics  . Remaining standalone components  orchestrate the execution of others. Components are also grouped into subsystems, with their background colored according to the teams responsible for their development. Numbers  , roman numerals  and letters  are used as labels throughout section 3.
Data Collection.
To fulfill the requirements described in section 2, the Ocean Guard tool required comprehensive data acquisition capabilities. The Data Acquisition subsystem collects data from multiple heterogeneous sources: 3rd party Providers (1), Physical Sensors (2), and Data Crawlers (3). The data is stored in one of the two data stores: the Label Store (A) for known data about anomalies – which may come from the 3rd Party Providers (1) or the Web Application (1) – and theRaw Data Store (B) for unprocessed data.
Continuous Training.
The next step in the implementation was to enable continuous training. Four components were developed for this purpose: a Synthetic Data Generation Pipeline (I); a Data Augmentation Pipeline (II); a Rule-Based Training Pipeline (III); and a ML-Based Training Pipeline (IV). The creation of these pipelines is an opportunity to introduce two critical model management tools: the Metadata Store (F) and the Model Registry (G), which together facilitate model lineage tracking and versioning.
Serving.
The Batch Prediction Pipeline (VIII) facilitates efficient processing of large-scale offline predictions, while results are stored in a dedicated Prediction Store (H). Complementing this offline processing capability, the API Prediction Service (8) is the main component of the system, providing low-latency inference capabilities to support real-time application requirements.
Monitoring.
The Ocean Guard tool requires a Governance Application (7) to enable model monitoring and analysis of system usage. This application interfaces with a dedicated Telemetry Store (I) that will store data about the use of the system while it runs.
Development.
Currently, software engineers, data scientists, and machine learning engineers can access the Ocean Guard source code maintained in the centralized Code Repository (C). Developers can easily modify the source code by employing one of the pre-configured tools: Integrated Development Environments (IDEs) (4) for structured software engineering tasks, and Notebooks (5) for interactive prototyping and analysis. This approach accommodates the diverse workflows and preferences of a multidisciplinary team.
Crontinuous Delivery.
Following modern software engineering practices, a Continuous Integration (CI) Pipeline (v) is used to build and test code. After that, dedicated Continuous Delivery (CD) Pipelines (VI) deploy traditional software-based components, whereas specialized Continuous Delivery of Machine Learning (CD4ML) Pipelines (VII) do the same for ML-based components. These pipelines interact with a centralized Artifact Registry (D) that maintains versioned software components, containerized applications and serialized models, ensuring reproducibility and traceability throughout the development cycle.
4 API Architecture
This section dives deeper into the architecture of the Ocean Guard API, which corresponds to the API Prediction Service (8) in the architecture presented in section 3. The Ocean Guard API is the center of the ML-Enabled System, since it applies the rule-based and machine-learning based anomaly detection models over the data received by the system.
Figure 2: Component Architecture of the Ocean Guard API. It follows the hexagonal architecture [12]. The dark blue hexagon represents the Business Logic, where Domain Driven Design patterns such as Entities, Value Objects and Services [4] are implemented. The light blue hexagon represents the Dependencies, where the Ports and Adapters pattern is implemented as an Anti-Corruption Layer against external dependencies [12]. Rectangles illustrate the external dependencies themselves, with an orange rectangle representing a Reading Dependency and and yellow rectangle representing an Output Dependency. Lastly, circles represent the different Ports and Adapters themselves, one per dependency. Solid arrows, going from a port to an adapter, represent the inside-out dependency flow, going from the Core to Ports then Adapters.
Figure 2 represents the key components of the API according to a hexagonal architecture, also known as onion or clean architecture [12]. This architectural pattern is designed to enhance flexibility and maintainability of a system by ensuring that its core business logic remains independent of external technologies.
According to the hexagonal architecture pattern, the system can be divided into three key elements: core, ports and adapters. The system follows a dependency rule: As one moves inward towards the core, the level of abstraction and the scope of policy increases. Specifically, the deeper layers of the architecture encapsulate higher-level policies and overarching business rules, while the outer layers handle more concrete implementation details. The core, being the innermost layer, represents the most general and highest level of abstraction [12].
Core.
Responsible for implementing the business logic. The Ocean Guard API follows the patterns of Domain Driven Design (DDD) [4]: entities and value objects represent the key objects related to the maritime domain. The services, also known as use cases [12], orchestrate the lifecycle of these objects.
Ports.
Responsible for establishing the contract and managing the communication between the core and the adapters. Within the Ocean Guard API, key components such as database repositories, dependency injection, security mechanisms, and web routers are defined as ports, facilitating the necessary integrations and ensuring that external concerns do not directly impact the core business logic.
Adapters.
Responsible for communication with external dependencies, making key details of the business logic. Adapters are categorized based on their dependencies: reading and output. Within the Ocean Guard API, reading adapters include the model adapter, third-party API adapter, storage adapter, database adapter, and configuration adapter, whereas output adapters include the web adapter and the cache adapter.
5 Team Configuration and Development Workflow
As a ML-enabled system, developing the Ocean Guard tool requires multiple specialists, including software engineers, data scientists, and machine learning engineers [10]. Table 1 summarizes the project organizational structure, which consists of four teams working in parallel, yet collaborative streams.
Table 1: Ocean Guard Teams. The Research and Innovation teams focus on different AI-based and algorithmic techniques. Whilst the work from them is research- and innovation-oriented, their work directly contributes to the Ocean Guard API Model Adapter. The Core Development team focuses on the API whereas the UI Development focuses on the WebApp.
The project workflow is deliberately structured to minimize interdependency between teams. Each team operates within clearly defined boundaries whose goal is to allow parallel development without blocking dependencies. The integration points happen via the CI/CD pipeline [5], where independently developed components can be tested against predefined contracts before merging into the main development branch [16]. This design technique is further explored in section 6.
Research Team.
It consists of Master students completing their theses. This team primarily focuses on exploring state-of-the-art techniques in the maritime domain. Although primarily research-oriented, their work directly contributes data products to the API, particularly to the API Model Adapter.
Innovation Team.
It consists of technical trainees focused on the exploration and implementation of state-of-the-practice techniques. This team employs rapid prototyping methodologies within constrained timelines to deliver demonstrable improvements to the API Model Adapter.
Core Development Team.
It consists of the authors of this paper focused on the fundamental aspects of the project infrastructure, including back-end development and infrastructure management. This team focuses on the API, databases, and model repository, with particular emphasis on refining the central API that serves as the integration point for all microservices.
UI Development.
It consist of hired web developers specializing in user interface design and front-end development. This team focuses primarily on the WebApp, which will be used by the end users for maritime tracking activities.
As the Research and Innovation teams consist of students and trainees, they operate under specific academic deadlines. These external constraints have been factored into the development process, further reinforcing the decision to implement a parallel work approach.
6 Contract-based development
As an ML-enabled system with a microservice architecture, the Ocean Guard tool was a complex system, as discussed in section 3. A recommended modeling technique for systems with such an architecture is Design by Contract [7]: an approach that establishes clear interfaces and agreements between software components or teams, defining how they interact without requiring knowledge of each other’s internal implementation [7].
In the development of the Ocean Guard tool, contract-based development was systematically implemented across the boundaries of interaction of the four teams introduced at section 5. These contracts served as formal specifications defining interaction patterns without constraining the internal implementations [13].
6.1 Code Contracts
For this project, Code Contracts document the expected behavior between two services – owned by different teams – which interacting synchronously or asynchronously via the network, usually under the HTTP protocol. As such, these specifications include endpoints, request parameters, and expected response.
There are two main code contracts in the Ocean Guard tool:
  • Between Data Crawlers (3) and external data sources. Creating these contracts provides the Core Development team with an anti-corruption layer [4] against the external data sources. This allowing them to create tests that verify the expected behavior of 3rd party APIs, which may change without notice.
  • Between the API Prediction Service (8) and the Web Application (8). Creating this contract allows the UI Development team to mock the responses of the API Prediction Service (8). This allows them to develop the front-end in parallel with the back-end developed by the Core Development team.
6.2 Data Contracts
For this project, Data Contracts document the expected formats in a data storage whose data is produced and consumed by services or pipelines owned by different teams. As such, these specifications include types, formats, distributions (when applicable or known) of data, besides the protocol for reading and writing the data itself.
There are three main data contracts in the Ocean Guard tool:
  • At the Label Store (A), between the 3rd Party Applications (1) and Web Application (9) that produce them, and the Notebooks (5), Synthetic Data Generation Pipeline (I) and Data Augmentation Pipeline (II) that consume them. These contracts ensure the Research and Innovation teams know what to expect to process labels.
  • At the Raw Data Store (B), between the 3rd Party Applications (1), Physical Sensors (2) and Data Crawlers (3) that produce them, and the Notebooks (5), Synthetic Data Generation Pipeline (I), Data Augmentation Pipeline (II) and Batch Prediction Pipeline (VIII) that consume them. These contracts ensure the Research and Innovation teams know what to expect to process raw data, as well as what their models might use during prediction.
  • At the Data Store (E), between the Synthetic Data Generation Pipeline (I) and Data Augmentation Pipeline (II) that produce them, and the Rule-Based Training Pipeline (III) and ML-Based Training Pipeline (IV) pipelines that consume them. These contracts ensure the Research and Innovation teams know how to consume the post-processed raw data and labels created by each other.
6.3 Model Contracts
For this project, Model Contracts document the expected input and output of a rule-based or ML-based model, as well as the format on which a trained model should be stored. As such, these specifications include data types, file formats, and data scientist–defined hyperparameters.
There is one main model contract in the Ocean Guard tool, at the Model Store (B), between the Rule-Based Training Pipeline (III) and ML-Based Training Pipeline (IV) pipelines that create them, and both the Batch Prediction Pipeline (VIII) and API Prediction Service (9) that use them. These contracts ensure the Research and Innovation teams know what to create for the Core Development team to use in the Ocean Guard tool back-end.
7 Challenges
In the experience of the authors, there were three main challenges for the development of the Ocean Guard tool: coupling, alignment, and communication.
Coupling.
As an ML-enabled system, the Ocean Guard tool was intrinsically complex, as discussed in section 3. Any modification to one of its components – either on code, model, or data – can quickly cascaded through the system. Therefore, mitigating the drawbacks of coupling was an important architectural challenge, particularly since services were being implemented concurrently during development.
Alignment.
As an ML-enabled system, the Ocean Guard tool required specialized developers, as discussed in section 5. Four teams worked simultaneously on different components, integrated via the Ocean Guard API. Therefore, keeping all team members oriented towards the same goal was an important organizational challenge, particularly due to the time-consuming nature of coordination activities.
Communication.
As an ML-enabled system with a microservices architecture, the Ocean Guard tool was a complex system, as discussed in section 3. Therefore, communicating the evolution of the Ocean Guard tool to internal and external stakeholders – with diverse technical backgrounds – was an important project management challenge, particularly since there were frequent feedback sessions due to its agile development process.
8 Lessons Learned
The challenges described in section 7 led to two lessons learned, related to applying techniques from the microservices literature that can be suitable for ML-enabled system development: contract-based design and ubiquitous language.
Contract-Based Design.
As previously mentioned in section 6, the authors employed design by contract [7] during the development of the Ocean Guard tool, a recommended modeling technique for systems with a microservices architecture [16]. This approach facilitated the definition of clear and precise specifications between components, setting expectations for interfaces and behaviors.
By applying design by contract through the development process, it tackled two challenges above: coupling, by reducing potential integration issues and improving system cohesion, and alignment, making it easier for the team to discuss how to integrate the components of the Ocean Guard tool.
Ubiquitous Language.
As previously mentioned in section 3, the authors employed Domain-Driven Design (DDD) [4] during the development of the Ocean Guard tool, a recommended modeling technique for systems with a microservices architecture [14]. This approach facilitated the creation of a shared vocabulary between developers and stakeholders, grounding discussions about the evolution of the system.
By refining a ubiquitous language through the development process, it tackled two challenges above: communication, by deepening stakeholders’ and developers’ understanding of the system and improving delivery feedback, and alignment, making it easier for the team to propose how to evolve the components of the Ocean Guard tool.
9 Conclusions
As the digital transformation of the maritime industry has accelerated, ships now function as floating data centers equipped with hundreds of sensors.
To extract actionable insights from this vast amount of data, this paper describes the Ocean Guard tool: a Machine Learning-Enabled System (MLES) for anomaly detection in the maritime domain. Its microservices architecture enables multiple specialized teams to work on the project in parallel, facilitated by a contract-based design approach.
Developing the Ocean Guard tool presented three major challenges:
  • coupling, related to cascading changes throughout the system;
  • alignment, related to coordinating four specialized teams working concurrently); and
  • communication, related to explaining the evolving system to stakeholders with diverse technical backgrounds.
To address them, the authors applied two techniques from the microservices literature during development:
  • contract-based design, which established clear specifications between components and reduced integration issues; and
  • ubiquitous language, which created a shared vocabulary among all participants via the application of Domain Driven Design (DDD).
Contract-based design effectively mitigated coupling between components and improved alignment among the team, whereas the ubiquitous language enhanced communication with stakeholders and further strengthened alignment among the team. Therefore, this case study in the maritime domain illustrates how established software engineering practices can be successfully adapted to complex Machine Learning-Enabled Systems.
Future work will focus on continuing the development of the Ocean Guard tool to ensure all functional and non-functional requirements are met. The development process will remain guided by the established contract-based approach and shared ubiquitous language. Moreover, the developers will continue to explore and implement both state-of-the-art and state-of-the-practice techniques with the Research and Innovation teams. These parallel exploration streams will enable the Ocean Guard tool to evolve with the latest advancements in maritime anomaly detection.